P1,000,000.00 pesos is needed at the end of each year to maintain each building in Rizal Technological University in Mandaluyong. If money increases at 12% per annum in a investment, how much money need to invest directly so the local government of Mandaluyong can reduce the annual cost by P250,000? *



1
Expert's answer
2021-11-23T11:02:54-0500

Solution:

Present value = "\frac{FV}{(1 + r)^{n} }"

FV = 750,000

r = 12% = 0.12

n = 1

Present value = "\frac{750,000}{(1 + 0.12)^{1} } = \frac{750,000}{1.12 } = 669,642.86"


Amount required to be invested directly to reduce the annual cost = P 669,642.86